40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2654     Entry time: Thu Mar 4 02:25:14 2010
Author: Jenne 
Type: Update 
Category: COC 
Subject: Further details on the magnet story, and SRM guiderod glued 

[Koji, Jenne]

First, the easy story:  SRM got it's guiderod & standoff glued on this evening.  It will be ready for magnets (assuming everything is sorted out....see below) as early as tomorrow.  We can also begin to glue PRM guiderods as early as tomorrow.

The magnet story is not as short.....

Problem: ITMX and ITMY's side magnets are not glued in the correct places along the z-axis of the optic (z-axis as in beam propagation direction). 

ITMX (as reported the other day) has the side magnet placement off by ~2mm.  ITMX side was glued using the magnet fixture from MIT and the teflon pads that Kiwamu and I improvised.

It was determined that the improvised teflon pads were too thin (maybe about 1m thick), so I took those out, and replaced them with the teflon pads stolen from the 40m's magnet gluing fixture.   (The teflon pad from the MIT fixture and the ones from the MIT fixture are the same within my measuring ability using a flat surface and feeling for a step between them.  I haven't yet measured with calipers the MIT pad thickness).  The pads from the 40m fixture, which were used in the MIT fixture to glue ITMY side last night were measured to be ~1.7mm thick.

Today when Koji hung ITMY, he discovered that the side magnet is off by ~1mm.  This improvement is consistent with the switching of the teflon pads to the ones from the 40m fixture.

We compared the 40m fixture with the one from MIT, and it looks like the distance from the edge of where the optic should sit to the center of the hole for the side magnet is different by ~1.1mm.  This explains the remaining ~1mm that ITMY is off by. 

We should put the teflon pads back into the 40m fixture, and only use that one from now on, unless we find an easy way to make thicker teflon pads for the fixture we received from MIT.  (The pads that are in there are about the maximum thickness that will fit).  I'm going to use my thickness measurements of SRM (taken in the process of gluing the guiderods) to see what thickness of pads / what fixture we want to actually use, but I'm sure that the fixture we found in the 40m is correct.  We can't use this fixture however, until we get some clean 1/4-28 screws.  I've emailed Steve and Bob, so hopefully they'll have something for us by ~lunchtime tomorrow. 

The ITMX side magnet is so far off in the Z-direction that we'll have to remove it and reglue it in the correct position in order for the shadow sensor to do anything.  For ITMY, we'll check it out tomorrow, whether the magnet is in the LED beam at all or not.  If it's not blocking the LED beam enough, we'll have to remove and reglue it too. 

Why someone made 2 almost identical fixtures, with a 1mm height difference and different threads for the set screws, I don't know.  But I don't think whoever that person was can be my friend this week. 

ELOG V3.1.3-